Can someone please help me solve this problem? Let f(x)=2x/x-1. Find the inverse function f^-1(x).
Answer by venugopalramana(3286)

Can someone please help me solve this problem? Let f(x)=2x/x-1. Find the inverse function f^-1(x).
Let f(x)=2x/x-1 = Y SAY...WE HAVE TO FIND X INTERMS OF Y
2X=YX-Y
Y=YX+2X=X(Y+2)
X = Y/(Y+2)
HENCE F^-1(X)=F(X)/{F(X)+2}